# Weighted shift operators on spaces of analytic functions over an annulus - I

### v1

@@ 4. Adjoints of unilateral weighted shifts
To see that (iii) implies (i), suppose that $\sum_{n=1}^{\infty}\left\lvert\frac{a_{n}}{w_{0}w_{1}\cdots w_{n-1}}\right\rvert^{p}<\infty$ holds. We apply the chaoticity criterion to show that $F_{w}^{*}$ is chaotic on $(\ell^{p}_{a,b}(\mathbb{D}))^{*}.$ Consider $D=\text{span}\{k_{n}(z):n\geq 0\}$. Then by the previous Proposition, the set $D$ is dense in $(\ell^{p}_{a,b}(\mathbb{D}))^{*}.$ Define $S:D\rightarrow D$ by $S(k_{n})=\frac{1}{w_{n}}k_{n+1},~{}~{}\forall~{}~{}n\geq 0.$ Clearly, $F_{w}^{*}S(f)=f,$ $\forall$ $f\in D,$ and $\sum_{n=0}^{\infty}(F_{w}^{*})^{n}(f)$ converges unconditionally for each $f\in D.$ We just need to show that $\sum_{n=0}^{\infty}S^{n}(f)$ converges unconditionally, for each $f\in D.$ We prove that $\sum_{n=1}^{\infty}\frac{1}{w_{0}w_{1}\cdots w_{n-1}}k_{n}$ is unconditionally convergent in $\ell^{p}_{a,b}(\mathbb{D})^{*}.$ Since $k_{n}=a_{n}f_{n}^{*}+b_{n-1}f_{n-1}^{*}$, for all $n\geq 1$ and $k_{0}=a_{0}f_{0}^{*}.$ Therefore, we have
@@
for all $n\geq N$. The right hand side of the above inequality is the $\ell^{p}$ element $\left\{\frac{a_{n+1}}{w_{0}\cdots w_{n}}\right\}_{n=1}^{\infty},$ and hence it is absolutely $p$-summable. Consequently, the series $\sum_{n\geq N}\frac{1}{w_{0}w_{1}\cdots w_{n-1}}k_{n}$ is convergent. The unconditional convergence occurs because $\{f_{n}^{*}\}$ is an unconditional Schauder basis. Hence, $F_{w}^{*}$ satisfies the chaoticity criterion, and (i) follows.
@@
($F_{w}=$ nuclear $+$ a standard weighted shift) Choose $a_{0}=b_{0}=w_{0}=1,$ and
$a_{n}=\frac{1}{n2^{n-1}},~{}b_{n}=\frac{1}{(n+1)2^{2n}},~{}w_{n}=4,~{}n\geq 1.$
Then $F_{w}$ is bounded on $\ell^{p}_{a,b}(\mathbb{D}),$ as it satisfies the conditions of Theorem 4.1. By Theorems 4.4 and 4.6, the adjoint $F_{w}^{*}$ is mixing and chaotic on $\ell^{p}_{a,b}(\mathbb{D})^{*}.$
We close this section with a remark. From the matrix representation of $F_{w}$, it follows that $F_{w}$ is similar to an operator of the form $A+F_{\alpha}$ for some bounded operator $A$ and a weighted forward shift $F_{\alpha}$ on $\ell^{p}(\mathbb{N})$. Hence, the adjoint $F_{w}^{*}$ is similar to a perturbation of a weighted backward shift on $\ell^{q}(\mathbb{N})$.
@@ 5. Notable differences: Eigenvectors, periodic vectors, and a zero-one law
We show that the forward weighted shifts $F_{w}$ on our spaces $\ell^{p}_{a,b}(\mathtt{A})$ and $\ell^{p}_{a,b}(\mathbb{D})$ are in general very different from the classical bilateral/unilateral weighted shifts. Below, we show in both the cases, that the adjoint of an unweighted forward shift operator can have a non-zero periodic vector without being hypercyclic. It is well known that if a bilateral weighted shift on $\ell^{p}(\mathbb{Z})$ has an orbit having a non-zero limit point, then the operator has a dense orbit, cf. [12] and [11]. Such a property is called as a zero-one law of orbital limit points for an operator. Other distinguishing properties are stated in this section.
@@
Let $|\lambda|=1$. Consider the space $\ell^{p}_{a,b}(\mathbb{D})$, $1<p<\infty$, where
$a_{n}=1~{}\text{and}~{}b_{n}=\frac{-1}{\lambda}\frac{n+1}{n+2},~{}n\geq 0.$
Then the unweighted forward shift $F$ is bounded, non-hypercyclic, and it is similar to $S+N$, where $S$ is the right shift on $\ell^{p}$ and $N$ is some nuclear (hence, compact) operator. Further,
(i) if $\lambda$ is a root of the unity, then $F^{*}$ has a non-trivial periodic vector in $\ell^{p}_{a,b}(\mathbb{D})^{*}$.
(ii) if $\lambda$ is not a root of the unity, then the zero-one law of orbital limit points fails for $F^{*}$.
@@
By Theorem $4.1$ we see that $F$ is bounded. However, to get an exact picture of $F$, we recall the matrix of $F$:
$[F]:=\begin{bmatrix}0&0&0\cdots\\ 1&0&0\ddots\\ \frac{1}{\lambda}\frac{1}{2.3}&1&0\ddots\\ \frac{1}{\lambda^{2}}\frac{1}{2.4}&\frac{1}{\lambda}\frac{1}{3.4}&1\ddots\\ \frac{1}{\lambda^{3}}\frac{1}{2.5}&\frac{1}{\lambda^{2}}\frac{1}{3.5}&\frac{1}{\lambda}\frac{1}{4.5}\ddots\\ \frac{1}{\lambda^{4}}\frac{1}{2.6}&\frac{1}{\lambda^{3}}\frac{1}{3.6}&\frac{1}{\lambda^{2}}\frac{1}{4.6}\ddots\\ \vdots&\ddots&\ddots~{}~{}~{}\ddots\end{bmatrix}.$
Now by the Theorem 4.1 we see that $F$ is bounded on $\ell^{p}_{a,b}(\mathbb{D}).$ Going back to the proof of Theorem $4.1$, we see that each $F_{n}$ is nuclear and $\|F\|_{n}\leq\frac{1}{n^{2}}$, $n\geq 1$. Hence, the sum $\sum_{n\geq 1}F_{n}$ is convergent in with respect to the nuclear norm. Consequently, $F$ is similar to “nuclear + right shift” on $\ell^{p}(\mathbb{N})$.
Again by Theorem 4.4 we know that $F^{*}$ is hypercyclic on $\ell^{p}_{a,b}(\mathbb{D})^{*}$ if and only if
$\lim\inf_{n\to\infty}(\lvert a_{n}\rvert+\lvert b_{n-1}\rvert)=0.$
However, in our case this condition is not satisfied. Hence $F^{*}$ is not hypercyclic.
The forward shift operator $F$ in $\ell^{p}_{a,b}(\mathbb{D})$ is given by $F(f(z))=zf(z),$ for all $f\in\ell^{p}_{a,b}(\mathbb{D})$ and $z\in\mathbb{D}.$ Now if $F$ is bounded on $\ell^{p}_{a,b}(\mathbb{D}),$ then for $f\in\ell^{p}_{a,b}(\mathbb{D})$ and $w\in\mathbb{D},$ we have
@@
that is, $F^{*}(ev_{z})=z(ev_{z}).$ It shows that $z$ is an eigenvalue of $F^{*}$ corresponding to the eigenvector $ev_{z}$ on $\ell^{p}_{a,b}(\mathbb{D})^{*}.$ This is in fact a well known fact, cf. Godefroy and Shapiro [20]. When $a_{n}=1$ and $b_{n}=\frac{-1}{\lambda}\frac{n+1}{n+2},$ $n\geq 0,$ where $\lambda$ is a root of the unity. Now
$\sum_{n=0}^{\infty}\lvert f_{n}(\lambda)\rvert^{q}=\sum_{n=0}^{\infty}\left\lvert\left(1-\frac{1}{\lambda}\frac{n+1}{n+2}\lambda\right)\lambda^{n}\right\rvert^{q}<\infty,$
for $1<q<\infty.$ Therefore the natural domain of $\ell^{p}_{a,b}(\mathbb{D})$ is given by $\text{Dom}_{(p)}=\mathbb{D}\cup\{\lambda\}.$ Since $F^{*}(ev_{z})=\lambda(ev_{z}),$ for all $z\in\mathbb{D}.$ It follows that $ev_{\lambda}\in\ell^{p}_{a,b}(\mathbb{D})^{*}.$ Also, we have $F^{*}(ev_{\lambda})=\lambda(ev_{\lambda}).$
If $\lambda$ is a square root of unity, say $\lambda^{m}=1$, then $F^{*m}(ev_{\lambda})=ev_{\lambda}$. Hence $F^{*}$ has a nontrivial periodic vector, namely $ev_{\lambda}$, which gives $(i)$.
Now, assume that $\lambda$ is not a root of unity. It is well known, for such $\lambda$, that the sequence $\{\lambda^{n}:n\geq 1\}$ is dense in the unit circle of $\mathbb{C}$. As
$\{F^{*n}({ev}_{\lambda}):n\geq 0\}=\{\lambda^{n}ev_{\lambda}:n\geq 0\},$
it follows that the orbit of $ev_{\lambda}$ under $F^{*}$ has several non-zero limit points, and thus $(ii)$ holds. We remind that $F^{*}$ is not hypercyclic. ∎
@@
Consider the space $\ell^{p}_{a,b}(\mathtt{A})$, $1\leq p<\infty$, having a normalized Schauder basis consisting of $\{f_{n}(z)=(a_{n}+b_{n}z)z^{n}:n\in\mathbb{Z},\}$ where $a_{0}=1,$ $b_{0}=-\frac{1}{\lambda}\frac{1}{2}$and $a_{n}=1,$ $b_{n}=-\frac{1}{\lambda}\frac{n+1}{n+2},$ for all $n\geq 1,$ and $a_{n}=\frac{2^{n}}{2^{2}},$ $b_{n}=-\frac{1}{\lambda}2^{n},$ for all $n\leq-1.$ Then the following hold:
(i) If $\lambda$ is a root of the unity, then $F^{*}$ has a non-trivial periodic vector in $(\ell^{p}_{a,b}(\mathtt{A}))^{*}$, but it is not hypercyclic.
(ii) If $\lambda$ is not a root of the unity, then the zero-one law of orbital limit points fails for $F^{*}$.
@@
From the given values of $\{a_{n}\}_{n\in\mathbb{Z}}$ and $\{b_{n}\}_{n\in\mathbb{Z}},$ we get
$\sum_{n=-\infty}^{\infty}\lvert f_{n}(\lambda)\rvert^{q}=\sum_{n=-1}^{-\infty}\left\lvert\left(\frac{2^{n}}{2^{2}}-\frac{1}{\lambda}2^{n}\lambda\right)\lambda^{-n}\right\rvert^{q}+\sum_{n=0}^{\infty}\left\lvert\left(1-\frac{1}{\lambda}\frac{n+1}{n+2}\lambda\right)\lambda^{n}\right\rvert^{q}<\infty.$
for $1<q<\infty.$ Here $r=\frac{1}{2}$ and $R=1.$ Therefore the natural domain of $\ell^{p}_{a,b}(\mathtt{A})$ is given by $\text{Dom}_{(p)}=\mathtt{A}\cup\{\lambda\},$ where $\mathtt{A}$ is a annulus $\frac{1}{2}<\lvert z\rvert<1.$ Since $F^{*}(ev_{z})=\lambda(ev_{z}),$ for all $z\in\mathtt{A}.$ In-particular for $z=\lambda,$ that is for the extra point in the $\text{\text{Dom}}_{(p)}$ we get
$F^{*}(ev_{\lambda})=\lambda(ev_{\lambda}),$
and hence $F^{*}$ has a non-trivial periodic vector. Now let us show that $ev_{\lambda}\in(\ell^{p}_{a,b}(\mathtt{A}))^{*}.$ But this is cleared by Closed-graph theorem. For the given values of $\{a_{n}\}_{n\in\mathbb{Z}}$ and $\{b_{n}\}_{n\in\mathbb{Z}},$ we have
@@
Therefore, we get the matrix of $F$ corresponding to the given Schauder basis as
@@
Clearly, $\sum_{i=1}^{\infty}\lVert E_{-i}\rVert+\lVert D\rVert+\sum_{i=1}^{\infty}\lVert E_{i}\rVert<\infty.$ Therefore $F$ is bounded on $\ell^{p}_{a,b}(\mathbb{D}).$ Again by the Theorem 3.10 we know that $F^{*}$ is hypercyclic on $(\ell^{p}_{a,b}(\mathbb{D}))^{*}$ if and only if $\lim\inf_{n\to\infty}(\lvert a_{n}\rvert+\lvert b_{n-1}\rvert)=0$ and $\lim\inf_{n\to\infty}(\lvert a_{-n}\rvert+\lvert b_{-n-1}\rvert)=0.$ But, in our case, $\lim\inf_{n\to\infty}(\lvert a_{n}\rvert+\lvert b_{n-1}\rvert)\neq 0.$ Hence $F^{*}$ is not hypercyclic. ∎
@@
There exists a weighted forward shift $F_{w}$ on a space of the form $\ell^{p}_{a,b}(\mathbb{D})$, which is not similar to a forward weighted shift on $\ell^{p}$. A similar conclusion holds for bilateral forward shifts.

### v2

@@ 4. Adjoints of unilateral weighted shifts
Now we prove that (iii) implies (i), by applying the chaoticity criterion to show that $F_{w}^{*}$ is chaotic on $(\ell^{p}_{a,b}(\mathbb{D}))^{*}.$ Consider $D=\text{span}\{k_{n}:n\geq 0\}$ which is dense in $(\ell^{p}_{a,b}(\mathbb{D}))^{*}.$ Define $S:D\rightarrow D$ by $S(k_{n})=\frac{1}{w_{n}}k_{n+1},~{}~{}\forall~{}~{}n\geq 0.$ Clearly, $F_{w}^{*}S(f)=f,$ $\forall$ $f\in D,$ and $\sum_{n=0}^{\infty}(F_{w}^{*})^{n}(f)$ converges unconditionally for each $f\in D.$ We just need to show that $\sum_{n=0}^{\infty}S^{n}(f)$ converges unconditionally, for each $f\in D.$ It suffices to prove that $\sum_{n=1}^{\infty}\frac{1}{w_{0}w_{1}\cdots w_{n-1}}k_{n}$ is unconditionally convergent in $(\ell^{p}_{a,b}(\mathbb{D}))^{*}.$ Recall that $k_{n}=a_{n}f_{n}^{*}+b_{n-1}f_{n-1}^{*}$, for all $n\geq 1$ and $k_{0}=a_{0}f_{0}^{*},$ and so we have
@@
for all $n\geq N$. The right hand side of the above inequality is the $\ell^{q}$ element
$\left\{\frac{a_{n+1}}{w_{0}\cdots w_{n}}\right\}_{n=1}^{\infty},$
and hence it is absolutely $q$-summable. Consequently, the series $\sum_{n\geq N}\frac{1}{w_{0}w_{1}\cdots w_{n-1}}k_{n}$ is convergent. The unconditional convergence occurs because $\{f_{n}^{*}\}$ is an unconditional Schauder basis. Hence, $F_{w}^{*}$ satisfies the chaoticity criterion, and (i) follows.
@@
Choose $a_{0}=b_{0}=w_{0}=1,$ and $a_{n}=\frac{1}{n2^{n-1}},~{}b_{n}=\frac{1}{(n+1)2^{2n}},~{}w_{n}=4,~{}n\geq 1$. Then $F_{w}$ is bounded on $\ell^{p}_{a,b}(\mathbb{D}),$ as it satisfies the conditions of Theorem 4.1. By Theorems 4.4 and 4.6, the adjoint $F_{w}^{*}$ is mixing and chaotic on $(\ell^{p}_{a,b}(\mathbb{D}))^{*},$ $1<p<\infty$.
A more general result on the dynamics of $F_{w}^{*}$ is given as follows in the next theorem. We omit its proof. A bilateral analogue was stated in Theorem 3.13.
Theorem 4.8.
Let $\mathcal{E}$ be a Banach space of analytic functions on the unit disc $\mathbb{D}$, having norm-separable dual $\mathcal{E}^{*}$. Let $k_{n}\neq 0$ for all $n\geq 0$. The following hold, assuming that the operator $F_{w}$ is bounded on $\mathcal{E}$.
  • (i)
    If $\inf_{n\geq 0}\big{(}|\prod_{k=1}^{n}w_{k}|^{-1}\|k_{n+1}\|\big{)}=0$, then $F_{w}^{*}$ is hypercyclic.
  • (ii)
    If $\lim_{n\rightarrow\infty}\big{(}|\prod_{k=1}^{n}w_{k}|^{-1}\|k_{n+1}\|\big{)}=0$, then $F_{w}^{*}$ is mixing.
  • (iii)
    If $\mathcal{E}$ is reflexive, and
    $\sum_{n\geq 1}\left|\left(\prod_{k=1}^{n}w_{k}\right)^{-1}\frac{f^{(n+1)}(0)}{{(n+1)}!}\right|<\infty,$
    for all $f\in\mathcal{E}$, then $F_{w}^{*}$ is chaotic.
  • (iv)
    The adjoint $F_{w}^{*}$ is supercyclic.
We close this section with results on compact perturbations, essential spectrum, and hypercyclic subspaces for $F_{w}^{*}$ when $F_{w}$ is defined on $\ell^{p}_{a,b}(\mathbb{D})$. Recall that the essential spectrum of a unilateral shift on $\ell^{2}$ was obtained in Shields [36]. The same proof works for $\ell^{p}$, $1<p<\infty$.
Proposition 4.9.
Suppose that the weighted shift $F_{w}$ is bounded on $\ell^{p}_{a,b}(\mathbb{D})$, where $1<p<\infty$. If $c_{n}:=w_{n+1}\frac{b_{n}}{a_{n+2}}-w_{n}\frac{a_{n}}{a_{n+1}}\frac{b_{n+1}}{a_{n+2}}\rightarrow 0$, as $n\rightarrow\infty$, then $F_{w}$ is similar to a perturbation $F_{\alpha}+K$ for some compact operator $K$ and a weighted forward shift $F_{\alpha}$ defined on $\ell^{p}$, having weights $\alpha_{i}=\frac{w_{i}a_{i}}{a_{i+1}}$, and hence $\sigma_{e}(F_{w})=\sigma_{e}(F_{\alpha})$.
Proof.
Recall from the matrix representation of $F_{w}$ on $\ell^{p}_{a,b}(\mathbb{D})$, it follows that $F_{w}$ is similar to an operator of the form $K+F_{\alpha}$ for some bounded operator $K$ and a weighted forward shift $F_{\alpha}$ on $\ell^{p}$. Here, $K=\sum_{n\geq 2}F_{n}$. Since $c_{n}$ is a null sequence, each $F_{n}$ becomes compact, and so $K$ is compact. As the essential spectrum is invariant under compact perturbations, we also have $\sigma_{e}(F_{w})=\sigma_{e}(F_{\alpha})$. ∎
Corollary 4.10.
Suppose that $F_{w}^{*}$ is hypercyclic on the dual $(\ell^{p}_{a,b}(\mathbb{D}))^{*}$, where $1<p<\infty$, and $11/p+1/q=$ 1. Under the assumptions as in the previous proposition, the adjoint $F_{w}^{*}$ possesses a hypercyclic subspace if and only if $F_{\alpha}^{*}$ has a hypercyclic subspace in $\ell^{q}(\mathbb{Z})$.
@@ 5. Notable differences: Eigenvectors, periodic vectors, and a zero-one law
We establish a couple of surprising results which show that the forward weighted shifts $F_{w}$ on the spaces $\ell^{p}_{a,b}(\mathtt{A})$ and $\ell^{p}_{a,b}(\mathbb{D})$ are very different from the classical bilateral/unilateral weighted shifts. The adjoint of an unweighted forward shift operator $F$ can have a non-zero periodic vector without $F$ being hypercyclic, which is in sharp contrast to the case of the classical weighted shifts. On another direction, a well known result states that if a bilateral weighted shift on $\ell^{p}(\mathbb{Z})$ has an orbit having a non-zero limit point, then the operator has a dense orbit, cf. [11]. (This property is known as a zero-one law of orbital limit points for an operator, also see [1], [8], and [10].) Such a law does not hold for $F_{w}$ defined on $\ell^{p}_{a,b}(\mathtt{A})$ and $\ell^{p}_{a,b}(\mathbb{D})$, as proved below. Further distinguishing characteristics are obtained in this section. These interesting properties of $F^{*}$ will be based on a suitable extension of the domain of the map $z\mapsto ev_{z}$, where $ev_{z}$ denotes the evaluation functional at $z$. The space we consider here, is based on Theorem $7$ in [2], where the authors studied the unweighted shift on a tridiagonal kernel space.
@@
Let $|\lambda|=1$. Consider a space of the form $\ell^{p}_{a,b}(\mathbb{D})$, $1<p<\infty$, wherein
$a_{n}=1,~{}0<|b_{n}|<1,~{}(n\geq 0).$
Suppose that the unweighted forward shift $F$ is bounded, and that
$\lim_{n\rightarrow+\infty}(b_{n}-b_{n+1})=0$.
Then, $F$ is non-hypercyclic, and it is similar to $S+K$, where $S$ is the right shift on $\ell^{p}$ and $K$ is some compact operator. Also, the following hold, assuming that
$\sum_{n\geq 0}|1+\lambda b_{n}|^{q}<\infty,$
$11/p+1/q=$ 1.
  1. (i)
    The zero-one law of orbital limit points fails for $F^{*}$.
  2. (ii)
    If $\lambda$ is a root of the unity, then $F^{*}$ has a non-trivial periodic vector in $(\ell^{p}_{a,b}(\mathbb{D}))^{*}$.
@@
By Theorem $4.1$ we can see that $F$ is a bounded operator, which is similar to a “compact + right shift” on $\ell^{p}(\mathbb{N})$.
By Theorem 4.4 (ii), we see that $F^{*}$ is not hypercyclic on $(\ell^{p}_{a,b}(\mathbb{D}))^{*}$.
The forward shift operator $F$ in $\ell^{p}_{a,b}(\mathbb{D})$ is given by $F(f(z))=zf(z),$ for all $f\in\ell^{p}_{a,b}(\mathbb{D})$ and $z\in\mathbb{D}.$ Now, since $F$ is bounded on $\ell^{p}_{a,b}(\mathbb{D}),$ for $f\in\ell^{p}_{a,b}(\mathbb{D})$ and $z\in\mathbb{D},$ we have
@@
that is, $F^{*}(ev_{z})=z(ev_{z}).$ It shows that $z\in\mathbb{D}$ is an eigenvalue of $F^{*}$ corresponding to the eigenvector $ev_{z}$ on $(\ell^{p}_{a,b}(\mathbb{D}))^{*},$ provided $ev_{z}$ is a non-zero functional. This fact is well known, cf. Godefroy and Shapiro [19].
In our case, $a_{n}=1,$ $0<\lvert b_{n}\rvert<1,$ $n\geq 0,$ and $\lambda$ is unimodular. Now,
$\sum_{n=0}^{\infty}\lvert f_{n}(\lambda)\rvert^{q}=\sum_{n=0}^{\infty}|1+\lambda b_{n}|^{q}<\infty,$ (5.1)
for $1<q<\infty.$ Therefore, the natural domain of $\ell^{p}_{a,b}(\mathbb{D})$ is given by $\text{Dom}_{(p)}=\mathbb{D}\cup\{\lambda\}.$ Note that if $f\in\ell^{p}_{a,b}(\mathbb{D})$, then $f(z)=\sum_{n\geq 0}\widehat{f}(n)f_{n}(z)$, and $\sum_{n\geq 0}|\widehat{f}(n)|^{p}<\infty$. In view of (5.1), we obtain that each $f(z)$ in $\ell^{p}_{a,b}(\mathbb{D})$ extends to $z=\lambda$, and thus, it is meaningful to define
$ev_{\lambda}(f)=f(\lambda):=\sum_{n\geq 0}\widehat{f}(n)f_{n}(\lambda),$
for all $f$. Since $|f(\lambda)|\leq M_{\lambda}\|f\|$ for some constant $M_{\lambda}>0$, we have that $ev_{\lambda}\in(\ell^{p}_{a,b}(\mathbb{D}))^{*}$. It follows that $F^{*}(ev_{\lambda})=\lambda(ev_{\lambda})$, i.e. $ev_{\lambda}$ is an eigenvector of $F^{*}$. Note that $ev_{\lambda}$ is non-zero as, for example, $ev_{\lambda}(g)\neq 0$, where
$g(z)=\sum_{n\geq 0}\frac{1}{\lambda^{n}2^{n}}\frac{|1+\lambda b_{n}|}{1+\lambda b_{n}}f_{n}(z)\in\ell^{p}_{a,b}(\mathbb{D}).$
If $\lambda$ is a root of unity, say $\lambda^{m}=1$, then $F^{*m}(ev_{\lambda})=ev_{\lambda}$. Hence, $F^{*}$ has a nontrivial periodic vector, namely $ev_{\lambda}$, which gives (ii).
To prove (i), first consider the orbit of $f:=ev_{z}+ev_{\lambda}$ under $F^{*}$, for any $z\in\mathbb{D}$. If $\lambda^{j}=1$, then $(F^{*})^{jn}(ev_{z}+ev_{\lambda})\rightarrow ev_{\lambda}$, as $n\rightarrow\infty$. In the other case when $\lambda$ is not a root of unity, we use a well known fact, for such $\lambda$, that the sequence $\{\lambda^{n}:n\geq 1\}$ is dense in the unit circle of $\mathbb{C}$, and obtain a sequence $(j_{n})$ such that $\lambda^{j_{n}}\rightarrow 0$, as $n\rightarrow\infty$. Consequently, we have $(F^{*})^{j_{n}}(ev_{z}+ev_{\lambda})\rightarrow ev_{\lambda}$, as $n\rightarrow\infty$. Thus, the zero-one aw of orbital limit points fails for $F^{*}$. We would like to recall that $F^{*}$ is not hypercyclic. ∎
Example 5.2.
For $|\lambda|=1$, if we set
$a_{n}=1~{}\text{and}~{}b_{n}=\frac{-1}{\lambda}\frac{n+1}{n+2},~{}n\geq 0,$
$F^{*}$ has the properties stated in (i) and (ii) of the above proposition.
@@
Let $|\lambda|=1$, and let $\mathtt{A}$ be an annulus $r<|z|<1$. Consider $\ell^{p}_{a,b}(\mathtt{A})$, $1<p<\infty$, where $10<|b_{n}|<$ 1 for all $n\in\mathbb{Z}$, and
$a_{n}=1,~{}n\geq 0,~{}~{}~{}~{}\text{and}~{}~{}~{}~{}0<|a_{n}|<1,~{}n\leq-1.$
If the right shift $F$ is bounded, and
$\sum_{n=2}^{\infty}\left\lvert\prod_{k=1}^{n-1}\frac{a_{-k}}{b_{-k-1}}\right\rvert^{p}<\infty,~{}~{}~{}~{}\sum_{n\geq 0}|1+\lambda b_{n}|^{q}<\infty~{}~{}~{}~{}\text{and}~{}~{}~{}~{}\sum_{n\geq 1}|a_{-n}+\lambda b_{-n}|^{q}<\infty,$
$11/p+1/q=$ 1, then the following hold.
  1. (i)
    The adjoint $F^{*}$ is non-hypercyclic.
  2. (ii)
    The zero-one law of orbital limit points fails for $F^{*}$.
  3. (iii)
    If $\lambda$ is a root of the unity, then $F^{*}$ has a non-trivial periodic vector in $(\ell^{p}_{a,b}(\mathtt{A}))^{*}$.
@@
In view of Proposition 3.15 (ii), we see that $F^{*}$ is not hypercyclic on $(\ell^{p}_{a,b}(\mathtt{A}))^{*}$. Indeed, using the basis expansion of $z^{n}$, $n\geq 1$, the set
$\{\|z^{n}\|:n\geq 1\}$
can be shown to be bounded.
Now,
$\sum_{n=-\infty}^{\infty}\lvert f_{n}(\lambda)\rvert^{q}=\sum_{n=1}^{\infty}|a_{-n}+\lambda b_{-n}|^{q}+\sum_{n=0}^{\infty}|1+\lambda b_{n}|^{q}<\infty,$ (5.2)
for $1<q<\infty.$ If $f\in\ell^{p}_{a,b}(\mathtt{A})$, then $f(z)=\sum_{n=-\infty}^{\infty}\widehat{f}(n)f_{n}(z)$, and $\sum_{n}|\widehat{f}(n)|^{p}<\infty$. In view of (5.2), the evaluation functional $ev_{\lambda}$ is well-defined and it is given by $ev_{\lambda}(f)=f(\lambda):=\sum_{n=-\infty}^{\infty}\widehat{f}(n)f_{n}(\lambda)$, for all $f$. Also, $ev_{\lambda}$ is continuous and non-zero functional because $ev_{\lambda}(g)\neq 0$, where
$g(z)=\sum_{n\geq 1}\frac{\lambda^{n}}{2^{n}}\frac{|a_{-n}+\lambda b_{-n}|}{a_{-n}+\lambda b_{-n}}f_{-n}(z)+\sum_{n\geq 0}\frac{1}{\lambda^{n}2^{n}}\frac{|1+\lambda b_{n}|}{1+\lambda b_{n}}f_{n}(z)\in~{}~{}\ell^{p}_{a,b}(\mathtt{A}).$
Hence, $ev_{\lambda}$ is an eigenvector of $F^{*}$. The remaining proof, similar to the unilateral case, yields $(ii)$ and $(iii)$. ∎
Example 5.4.
Let $\lvert\lambda\rvert=1,$ and consider
$a_{n}=1,~{}~{}~{}b_{n}=-\frac{-1}{\lambda}\frac{n+1}{n+2},~{}\forall n\geq 0,~{}~{}~{}\text{and}~{}~{}~{}a_{n}=\frac{2^{n}}{2^{2}},~{}~{}~{}b_{n}=-\frac{1}{\lambda}2^{n},~{}~{}~{}~{}\forall n\leq-1.$
Then, we have
@@
Therefore, we get the matrix of $F$ corresponding to $\{f_{n}\}$
@@
Note that the boundedness of $F$ cannot be conluded using Theorem 3.3. So, to show that $F$ is bounded, recall the matrix representation of $F$ and invoke the proof of Theorem 3.3. We see that $\sum_{i=1}^{\infty}\lVert E_{-i}\rVert+\sum_{i=1}^{\infty}\lVert E_{i}\rVert<\infty.$ Therefore, $F$ is bounded on $\ell^{p}_{a,b}(\mathtt{A}).$ For the given values of $\{a_{n}\}_{n\in\mathbb{Z}}$ and $\{b_{n}\}_{n\in\mathbb{Z}},$ we get
$\sum_{n=-\infty}^{\infty}\lvert f_{n}(\lambda)\rvert^{q}=\sum_{n=1}^{\infty}\left\lvert\left(\frac{1}{2^{2}2^{n}}-\frac{1}{\lambda 2^{n}}\lambda\right)\lambda^{-n}\right\rvert^{q}+\sum_{n=0}^{\infty}\left\lvert\left(1-\frac{1}{\lambda}\frac{n+1}{n+2}\lambda\right)\lambda^{n}\right\rvert^{q}<\infty,$ (5.4)
for $1<q<\infty.$ Here $r=\frac{1}{2}$ and $R=1.$ Therefore the natural domain of $\ell^{p}_{a,b}(\mathtt{A})$ is given by $\text{Dom}_{(p)}=\mathtt{A}\cup\{\lambda\},$ where $\mathtt{A}$ is a annulus $\frac{1}{2}<\lvert z\rvert<1.$ Then, $F^{*}$ has the properties stated in (i), (ii), and (iii) of the above proposition.
@@
There exists a weighted forward shift $F_{w}$ on a space of the form $\ell^{p}_{a,b}(\mathbb{D})$, which is not similar to a weighted shift on $\ell^{p}$. A similar conclusion holds for bilateral shifts as well.